Circular No. 6378

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) C. G. Mason, R. D. Gehrz, T. J. Jones, J. Mergen, and D. Williams,

University of Minnesota, report additional photometric magnitudes (cf. IAUC 6365) obtained on Apr. 10.1 UT, using a 27′′ circular aperture with a throw of 83′′ between the source and reference beams (uncertainties < 10 percent): [1.2 μm] = +4.45, [1.6 μm] = +4.26, [2.3 μm] = +4.06, [3.6 μm] = +1.62, [4.9 μm] = −0.04, [7.8 μm] = −2.93, [8.7 μm] = −3.36, [9.8 μm] = −3.84, [10.3 μm] = −3.88, [11.6 μm] = −4.31, [12.5 μm] = −4.16, [18 μm] = −4.26. Infrared spectral energy distributions show that the dust temperature is ∼ 20 percent higher than would be expected for small black spheres at the same heliocentric distance. The 10- and 20-μm silicate emission features are very weak.

A. T. Tokunaga, University of Hawaii; T. Y. Brooke, Jet Propulsion Laboratory; H. A. Weaver, Applied Research Corporation; and J. Crovisier and D. Bockelée-Morvan, Observatory of Paris, Meudon, report the detection with the NASA Infrared Telescope Facility of the P3 and P5 lines of the ν_{3} fundamental near 3.04 μm from C_{2}H_{2} (acetylene) emission on Apr. 8.2 UT. A very preliminary range for the production rate of acetylene is 0.5–1.0 × 10^{27} molecules/s (or 0.1–0.4 percent of the water abundance, assuming a water-production rate of 2.4–4.8 × 10^{29} molecules/s).

SUPERNOVA 1992ad IN NGC 4411B S. D. Van Dyk, University of California at Berkeley; R. A. Sramek,

National Radio Astronomy Observatory; K. W. Weiler and M. J. Montes, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute, ESA, report: “Observations of the type-II SN 1992ad made on 1992 Aug. 25.81 UT with the Very Large Array telescope show the first detection of radio emission at wavelength 3.6 cm (8.4 GHz) with a flux density of 0.294 ± 0.035 mJy at α = 12^{h}26^{m}49^{s}.593, δ = +8^{o}52′38′′.72 (equinox J2000.0; ± 0′′.2 in α, δ). This is coincident (to within 1′′) with the optical position given on IAUC 5552 (position end figures for equinox J2000.0: 49^{s}.62, 38′′.4). SN 1992ad was also observed at 3.6 cm on 1993 Feb. 2.49, at a flux density of 0.312 ± 0.057 mJy, and again at 3.6 cm on 1993 Oct. 25.69, when the source had risen in flux density to 0.787 ± 0.101 mJy. This behavior over the course of more than one year is consistent with the source’s being a radio supernova that is still optically thick to its radio emission. Radio monitoring is continuing at this and other wavelengths.”

1996 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green